OpenStudy (anonymous):

What did I do wrong: Solve the following simultaneous equation: y = x^2 and y = 5x - 6

OpenStudy (anonymous):

x^2 - 5x + 6 = 0 (x + 6) (x - 1) = 0 x + 6 = 0, x - 1 = 0 x = -6, x = 1, y = 36, y = 1 The answers say that the answers are x = 3, y = 9 and x = 2 and y = 4

OpenStudy (anonymous):

You factored wrong. The second line of your working is wrong. It should factor to (x-3)(x-2)=0

OpenStudy (anonymous):

How?

OpenStudy (anonymous):

You need two numbers that multiply to give +6 and add to give -5. So the factors of 6 are 6 and 1, or 2 and 3. Either both factors are positive, or both are negative (to get positive 6 as the 'c' in your quadratic equation. We choose 2 and 3, and make both of them negative because -2 + -3 = -5, which is the 'b' in your quadratic equation, i.e.the coefficient of the x term. Try expanding the second line of your working and see that you do not get back the original quadratic.

OpenStudy (anonymous):

Oh I remember that

OpenStudy (anonymous):

Now I am stuck on y = 2x^2 and y = 8x

OpenStudy (anonymous):

I thought the last number could be 4 but it doesn't add up to 8

OpenStudy (anonymous):

Make them equal to each other.

OpenStudy (anonymous):

2x^2 - 8x - _ = 0

OpenStudy (anonymous):

That is what I'm up to now

OpenStudy (anonymous):

Your last number, your c, is 0, right?

OpenStudy (anonymous):

So x is common to both 2x^2 and -8x, which means we can factor out just the x, which leaves us with x(2x-8)=0

OpenStudy (anonymous):

ok

OpenStudy (ajprincess):

u can also factor out 2 along vth x vch will giv u 2x(x-4)=0